OpenStudy (mic123):

the graph of g(x) is f(x) translated to the left 8 units and up 2 units what is the function rule for g(x) given f(x)=x^s?

OpenStudy (anonymous):

Okay, so \[f(x)=x ^{2}\] since it is translated left 8 units, then it would be x+8 in parenthesis. \[f(x)=(x+8)^{2}\] since it is translated up 2 units, the you would add two outside of the parenthesis \[f(x)=(x+8)^{2}+2\] now you replace f(x) with g(x) and that's you're new equation.\[g(x)=(x+8)^{2}+2\]